OpenStudy (newwar):

What is the solution to the inequality? 5(x−2)<−3x+6

OpenStudy (mathmale):

Distribute: 5(x-2). Your result?

OpenStudy (newwar):

2,10

OpenStudy (mathmale):

You need to mult. x by 5 and also to mult that -2 by 5. Your answer must contain that '-' sign. Try again. 5(x-2)=?

OpenStudy (angel.candy):

5x-10

OpenStudy (angel.candy):

That's what it is when you distribute. What do you think you do next?

OpenStudy (newwar):

do i simplify

OpenStudy (newwar):

wait never mind

OpenStudy (newwar):

so is that my equation

OpenStudy (angel.candy):

3 becomes - 3 The answer is -3 -3+6 evaluates to 3 The right hand side of your inequality can be written as: 3 5x-10 ‹ 3 Move the -10 to the right hand side by adding 10 to both sides, like this: To the left hand side: -10 + 10 = 0 The answer is 5x To the right hand side: 3 + 10 = 13 The answer is 13 Now, the inequality reads: 5x ‹ 13 To isolate the x, we have to divide both sides of the inequality by the other "stuff" (variables or coefficients) around the x on the left side of the inequality. he last step is to divide both sides of the inequality by 5 like this: To divide x by 1 The x just gets copied along in the numerator. The answer is x 5x ÷ 5 = x 13 ÷ 5 = The solution to your inequality is: x ‹ So, your solution is: 13/5 x must be less than

OpenStudy (newwar):

ok i get it now so it would be 13/5

OpenStudy (newwar):

thank you @Angel.Candy